With a wind behind it, Glenn's pigeon flies 72 miles from Santa Rosa to Cloverdale in 6 hours. The return trip takes 8 hours because the pigeon flies into the wind. If the wind speed is the same on each trip, find the speed of the pigeon in still air and find the wind speed.

I know how to solve a system of equations, but i dont know how to set this word problem up. For this problem, i am required to provide two variables and two equations. Please help.

Answer by mananth(16946)

pigeon speed =x mph
wind speed =y mph
against wind 8 hours
with wind 6 hours

Distance against 72 miles distance with 72 miles
t=d/r against wind -
72 / ( x - y )= 8
8 ( x - y ) = 72
8 x - 8 y = 72 ....................1

72 / ( x + y )= 6
6 ( x + y ) = 72
6 x + 6 y = 72 ...............2
Multiply (1) by 3
Multiply (2) by 4
we get
24 x + -24 y = 216
24 x + 24 y = 288
48 x = 504
/ 48
x = 10.50 mph

plug value of x in (1) y
8 x -8 y = 72
84 -8 -84 = 72
-8 y = 72
-8 y = -12 mph
y = 1.50
pigeon speed 10.50 mph
wind speed 1.50 mph

m.ananth@hotmail.ca